# Combating Confirmation Bias: A Unified Pseudo-Labeling Framework for Entity Alignment

### v2

@@ Abstract
Entity alignment (EA) aims at identifying equivalent entity pairs across different knowledge graphs (KGs) that refer to the same real-world identity. It has been a compelling but challenging task that requires the integration of heterogeneous information from different KGs to expand the knowledge coverage and enhance inference abilities. To circumvent the shortage of prior seed alignments provided for training, recent EA models utilize pseudo-labeling strategies to iteratively add unaligned entity pairs predicted with high confidence to the prior seed alignments for model retraining. However, the adverse impact of confirmation bias during pseudo-labeling has been largely overlooked, thus hindering entity alignment performance. To systematically combat confirmation bias, we propose a new Unified Pseudo-Labeling framework for Entity Alignment (UPL-EA) that explicitly alleviates pseudo-labeling errors to boost the performance of entity alignment. UPL-EA achieves this goal through two key innovations: (1) Optimal Transport (OT)-based pseudo-labeling uses discrete OT modeling as an effective means to determine entity correspondences and reduce erroneous matches across two KGs. An effective criterion is derived to infer pseudo-labeled alignments that satisfy one-to-one correspondences at each iteration; (2) Pseudo-label ensembling refines pseudo-labeled alignments by combining predictions over multiple pseudo-labeling iterations. The refined pseudo-labeled alignments are thereafter used to augment prior seed alignments to reinforce subsequent model training for alignment inference. The effectiveness of UPL-EA in eliminating pseudo-labeling errors is both theoretically supported and experimentally validated. Our extensive results and in-depth analyses demonstrate the superiority of UPL-EA over 15 competitive baselines and its utility as a general pseudo-labeling framework for entity alignment.
@@ 1 Introduction
Over the last decade, there has been a surge of research efforts dedicated to entity alignment across KGs. Most mainstream EA models are embedding-based; they embed KGs into a shared latent embedding space so that similarities between entities can be measured via their embeddings for alignment inference. To leverage structural information in KGs, more recent EA models exploit the power of Graph Neural Networks (GNNs) to encode KG structures for entity alignment. Methods like GCN-Align [51] utilize GCNs to learn better entity embeddings by aggregating features from neighboring entities. However, GCNs and their variants suffer from an over-smoothing issue [23, 13], where the embeddings of entities among local neighborhoods become indistinguishably similar as the number of convolution layers increases. To alleviate over-smoothing during GCN neighborhood aggregation, recent works [53, 52, 62] use a highway strategy [34] on GCN layers, which “mixes” the smoothed entity embeddings with the original features. Despite achieving competitive results, these methods require an abundant amount of pre-aligned entity pairs (known as prior seed alignments) provided for training, which are labor-intensive and costly to acquire in real-world KGs. To tackle the shortage of prior seed alignments, recently proposed models, such as BootEA [39], IPTransE [61], MRAEA [20], and RNM [62], adopt a bootstrapping strategy that iteratively selects unaligned entity pairs predicted with high confidence as pseudo-labeled alignments and adds them to prior seed alignments for model retraining. The bootstrapping strategy, originating from the field of statistics, is also referred to as pseudo-labeling—a predominant learning paradigm proposed to tackle label scarcity in semi-supervised learning.
@@
In general semi-supervised learning, pseudo-labeling approaches inherently suffer from confirmation bias [1, 44]. The confirmation bias refers to using incorrectly predicted labels generated on unlabeled data for subsequent training, thereby misleadingly increasing model confidence in incorrect predictions and leading to a biased model with degraded performance. Unfortunately, there is a lack of understanding of the fundamental factors that give rise to confirmation bias for pseudo-labeling-based entity alignment. Our analysis (see Section 2.2) advocates that the confirmation bias is exacerbated during pseudo-labeling for entity alignment. Due to the lack of sufficient prior seed alignments at the early stages of training, the existing models tend to learn uninformative entity embeddings and consequently generate error-prone pseudo-labeled alignments based on unreliable model predictions. We characterize pseudo-labeling errors into two types: (1) Type I pseudo-labeling errors refer to conflicted misalignments, where a single entity in one KG is simultaneously aligned with multiple entities in another KG with erroneous matches. (2) Type II pseudo-labeling errors refer to one-to-one misalignments, where an entity in one KG is incorrectly matched with an entity in another KG. The pseudo-labeling errors, if not properly mitigated, would propagate into subsequent model training, thereby jeopardizing the efficacy of pseudo-labeling-based entity alignment. However, current pseudo-labeling-based EA models have made only limited attempts to alleviate alignment conflicts using simple heuristics [61, 40, 20, 62] or imposing constraints to enforce hard alignments [39, 10], while the confirmation bias has been left under-explored.
@@
To address the research gap, we propose a novel unified pseudo-labeling framework for entity alignment (UPL-EA) aimed at alleviating confirmation bias and improving entity alignment performance. The key idea lies in “reliably” pseudo-labeling unaligned entity pairs based on model predictions and augmenting prior seed alignments to iteratively improve model performance. UPL-EA comprises two essential components: Optimal Transport (OT)-based pseudo-labeling and pseudo-label ensembling, to effectively reduce pseudo-labeling errors. OT-based pseudo-labeling considers entity alignment as a probabilistic matching process between entity sets in two KGs. An effective criterion is mathematically derived to select pseudo-labeled alignments that satisfy one-to-one correspondences, thus mitigating conflicted misalignments (Type I pseudo-labeling errors) at each iteration. Pseudo-label ensembling reduces the pseudo-label selection variability by finding consensus predictions over multiple pseudo-labeling iterations, thereby eliminating one-to-one misalignments (Type II pseudo-labeling errors). The refined pseudo-labeled alignments are used to augment prior seed alignments to reinforce subsequent model training for alignment inference. To our best knowledge, we are the first to address the confirmation bias inherent in pseudo-labeling-based entity alignment. Comprehensive experiments and analyses validate the superior performance of UPL-EA over state-of-the-art supervised and semi-supervised baselines and its utility as a general pseudo-labeling framework to improve entity alignment performance.
@@ 2.1 Problem Statement
A knowledge graph (KG) can be represented as $\mathcal{G}=\{\mathcal{E},\mathcal{R},\mathcal{T}\}$ with the entity set $\mathcal{E}$, relation set $\mathcal{R}$, and relational triplet set $\mathcal{T}$. Each triplet is denoted as $(e_{i},r,e_{j})\in\mathcal{T}$, which represents that a head entity $e_{i}\in\mathcal{E}$ is connected to a tail entity $e_{j}\in\mathcal{E}$ via a relation $r\in\mathcal{R}$. Each entity $e_{i}$ is characterized by an entity feature vector $\bm{x}_{i}\in\mathbb{R}^{f}$, which can be obtained from entity names with semantic meanings.
Formally, given two KGs, $\mathcal{G}_{1}=\{\mathcal{E}_{1},\mathcal{R}_{1},\mathcal{T}_{1}\}$ and $\mathcal{G}_{2}=\{\mathcal{E}_{2},\mathcal{R}_{2},\mathcal{T}_{2}\}$, the task of entity alignment (EA) aims to discover a set of one-to-one equivalent entity pairs =$\mathcal{I}=\{(e_{i},e_{j})\in\mathcal{E}_{1}\times\mathcal{E}_{2}|\;e_{i}\equiv e_{j}\}$ between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, where $e_{1}\in\mathcal{E}_{1}$, $e_{2}\in\mathcal{E}_{2}$, and $\equiv$ indicates an equivalence relationship between $e_{1}$ and $e_{2}$. In many cases, a small set of equivalent entity pairs $\mathcal{S}_{0}\in\mathcal{I}$, known as prior seed alignments, is provided beforehand and used as training data. Apart from entities included in $\mathcal{S}_{0}$, there are two sets of unaligned entities $\mathcal{E}^{U}_{1}\subset\mathcal{E}_{1}$ and $\mathcal{E}^{U}_{2}\subset\mathcal{E}_{2}$ in $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively.
In this work, we consider real-world scenarios where prior seed alignments $\mathcal{S}_{0}$ are usually scarce due to high labeling costs. We focus on the task of pseudo-labeling-based entity alignment, which aims to leverage both prior seed alignments and unaligned entity pairs to more effectively train an EA model in semi-supervised settings. This is achieved by iteratively selecting a set of unaligned entity pairs as pseudo-labeled alignments $\widehat{\mathcal{S}}\subset\mathcal{E}^{U}_{1}\times\mathcal{E}^{U}_{2}$ and using $\widehat{\mathcal{S}}$ to augment prior seed alignments $\mathcal{S}_{0}$, i.e., $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\widehat{\mathcal{S}}$, for subsequent model retraining.
@@ 2.2 Analysis of Confirmation Bias
The key to pseudo-labeling-based entity alignment lies in selecting reliable pseudo-labeled alignments to effectively boost model performance; otherwise, pseudo-labeling errors could propagate into subsequent model training, leading to confirmation bias [1]. To investigate the impact of confirmation bias on pseudo-labeling-based entity alignment, we perform an error analysis of a naive pseudo-labeling strategy used in previous studies [40]. This strategy simply selects pairs of unaligned entities whose embedding distances (defined using Eq. (6)) are smaller than a pre-specified threshold as pseudo-labeled alignments. Our analysis is carried out on a widely used cross-lingual KG pair, DBP15KZH_EN (with details presented in Section 4.1), using the conventional 30%-70% split ratio to randomly partition the 15,000 ground truth entity alignments into training and test data. To understand the underlying causes of confirmation bias, we explicitly calculate the numbers of Type I and Type II pseudo-labeling errors, along with the number of correctly pseudo-labeled alignments, at different pseudo-labeling iterations on the test data.
As shown in Fig. 1(a), the naive pseudo-labeling strategy selects only a relatively small number of correctly pseudo-labeled alignments in the first three iterations (although Type I and Type II pseudo-labeling errors seem low). Thus, the corresponding performance in Fig. 1(c) shows minimal improvements. From the fourth iteration, although the number of correctly pseudo-labeled alignments gradually increases, Type I and Type II pseudo-labeling errors also propagate and increase as the training proceeds. This accumulation of errors give rise to confirmation bias, severely hindering the capability of pseudo-labeling for performance improvements.
@@
Our analysis affirms that the confirmation bias essentially stems from Type I and Type II pseudo-labeling errors. These errors, if not adequately addressed, would propagate into subsequent model training, thus diminishing the effectiveness of pseudo-labeling for entity alignment. Our work is thus motivated to explicitly eliminate Type I and Type II pseudo-labeling errors. As evidenced in Fig. 1(b), our proposed Unified Pseudo-Labeling (UPL) strategy effectively eliminates all Type I pseudo-labeling errors across all iterations while mitigating the number of Type II pseudo-labeling errors, leading to substantial performance gains as shown in Fig. 1(c).
@@ 3 The Proposed UPL-EA Framework
With insights from our analysis in Section 2.2, the proposed UPL-EA framework is designed to systematically address confirmation bias for pseudo-labeling-based entity alignment. The core of UPL-EA is a new Unified Pseudo-Labeling (UPL) strategy that iteratively generates reliable pseudo-labeled alignments to enhance the entity alignment (EA) model. UPL utilizes two essential components: (1) OT-based pseudo-labeling, which generates pseudo-labeled alignments with one-to-one correspondences, effectively eliminating conflicted misalignments at each iteration (Addressing Type I pseudo-labeling errors); and (2) pseudo-label ensembling, which combines pseudo-labeled alignments generated through OT modeling over multiple iterations, reducing prediction variability and mitigating one-to-one misalignments across iterations (Addressing Type II pseudo-labeling errors).
Fig. 2 illustrates an overview of the proposed UPL-EA framework. At each iteration $t$, the EA model learns entity embeddings based on the available seed alignments $\mathcal{S}$. OT-Based Pseudo-Labeling then takes entity embeddings as input to generate pseudo-labeled alignments $\widehat{\mathcal{S}}_{t}$, which are combined with prior seed alignments $\mathcal{S}_{0}$ to train the EA model in the next iteration $t+1$. After every $m$ iterations, pseudo-labeled alignments generated over multiple iterations $\widehat{\mathcal{S}}_{t}|\;t=1,...,m\}$ are passed on to Pseudo-Label Ensembling for further refinement, preventing one-to-one misalignments from propagating into subsequent model training. The refined pseudo-labeled alignments $\widehat{\mathcal{S}}$ are used to augment prior seed alignments $\mathcal{S}_{0}$, i.e., $\mathcal{S}_{0}\leftarrow\mathcal{S}_{0}\cup\widehat{\mathcal{S}}$, which are subsequently used in the next $m$ iterations for model training. The EA model and the UPL strategy mutually reinforce each other to learn informative entity embeddings. Lastly, the learned entity embeddings are used for entity alignment inference.
Figure 2: An overview of the proposed UPL-EA framework.

### v3

@@ Abstract
Entity alignment (EA) aims at identifying equivalent entity pairs across different knowledge graphs (KGs) that refer to the same real-world identity. It has been a compelling but challenging task that requires the integration of heterogeneous information from different KGs to expand the knowledge coverage and enhance inference abilities. To circumvent the shortage of prior seed alignments provided for training, recent EA models utilize pseudo-labeling strategies to iteratively add unaligned entity pairs predicted with high confidence to the seed alignments for model training. However, the adverse impact of confirmation bias during pseudo-labeling has been largely overlooked, thus hindering entity alignment performance. To systematically combat confirmation bias, we propose a new Unified Pseudo-Labeling framework for Entity Alignment (UPL-EA) that explicitly alleviates pseudo-labeling errors to boost the performance of entity alignment. UPL-EA achieves this goal through two key innovations: (1) Optimal Transport (OT)-based pseudo-labeling uses discrete OT modeling as an effective means to determine entity correspondences and reduce erroneous matches across two KGs. An effective criterion is derived to infer pseudo-labeled alignments that satisfy one-to-one correspondences; (2) Parallel pseudo-label ensembling refines pseudo-labeled alignments by combining predictions over multiple models independently trained in parallel. The ensembled pseudo-labeled alignments are thereafter used to augment seed alignments to reinforce subsequent model training for alignment inference. The effectiveness of UPL-EA in eliminating pseudo-labeling errors is both theoretically supported and experimentally validated. Our extensive results and in-depth analyses demonstrate the superiority of UPL-EA over 15 competitive baselines and its utility as a general pseudo-labeling framework for entity alignment.
@@ 1 Introduction
Over the last decade, there has been a surge of research efforts dedicated to entity alignment across KGs. Most mainstream EA models are embedding-based; they embed KGs into a shared latent embedding space so that similarities between entities can be measured via their embeddings for alignment inference. To leverage structural information in KGs, more recent EA models exploit the power of Graph Neural Networks (GNNs) to encode KG structures for entity alignment. Methods like GCN-Align [52] utilize GCNs to learn better entity embeddings by aggregating features from neighboring entities. However, GCNs and their variants suffer from an over-smoothing issue [24, 13], where the embeddings of entities among local neighborhoods become indistinguishably similar as the number of convolution layers increases. To alleviate over-smoothing during GCN neighborhood aggregation, recent works [54, 53, 63] use a highway strategy [35] on GCN layers, which “mixes” the smoothed entity embeddings with the original features. Despite achieving competitive results, these methods require an abundant amount of pre-aligned entity pairs (known as prior seed alignments) provided for training, which are labor-intensive and costly to acquire in real-world KGs. To tackle the shortage of prior seed alignments, recently proposed models, such as BootEA [40], IPTransE [62], MRAEA [21], and RNM [63], adopt a bootstrapping strategy that iteratively selects unaligned entity pairs predicted with high confidence as pseudo-labeled alignments and adds them to seed alignments for subsequent model training. The bootstrapping strategy, originating from the field of statistics, is also referred to as pseudo-labeling—a predominant learning paradigm proposed to tackle label scarcity in semi-supervised learning.
@@
In general semi-supervised learning, pseudo-labeling approaches inherently suffer from confirmation bias [1, 45]. The confirmation bias refers to using incorrectly predicted labels generated on unlabeled data for subsequent training, thereby misleadingly increasing model confidence in incorrect predictions and leading to a biased model with degraded performance. Unfortunately, there is a lack of understanding of the fundamental factors that give rise to confirmation bias for pseudo-labeling-based entity alignment. Our analysis (see Section 2.2) advocates that the confirmation bias is exacerbated during pseudo-labeling for entity alignment. Due to the lack of sufficient seed alignments at the early stages of training, the existing models tend to learn uninformative entity embeddings and consequently generate error-prone pseudo-labeled alignments based on unreliable model predictions. We characterize pseudo-labeling errors into two types: (1) Conflicted misalignments, where a single entity in one KG is simultaneously aligned with more than one entity in another KG with erroneous matches, violating the one-to-one correspondence. (2) One-to-one misalignments, where an entity in one KG is aligned to a single but incorrect counterpart in another KG. The pseudo-labeling errors, if not properly mitigated, would propagate into subsequent model training, thereby jeopardizing the efficacy of pseudo-labeling-based entity alignment. However, current pseudo-labeling-based EA models have made only limited attempts to alleviate alignment conflicts using simple heuristics [62, 41, 21, 63] or imposing constraints to enforce hard alignments [40, 10], while the confirmation bias has been left under-explored.
@@
To address the research gap, we propose a novel Unified Pseudo-Labeling framework for Entity Alignment (UPL-EA) aimed at alleviating confirmation bias and improving entity alignment performance. The key idea lies in “reliably” pseudo-labeling unaligned entity pairs based on model predictions and augmenting seed alignments to iteratively improve model performance. UPL-EA comprises two essential components: Optimal Transport (OT)-based pseudo-labeling and parallel pseudo-label ensembling, to effectively reduce pseudo-labeling errors. OT-based pseudo-labeling considers entity alignment as a probabilistic matching process between entity sets in two KGs. An effective criterion is mathematically derived to select pseudo-labeled alignments that satisfy one-to-one correspondences, thus mitigating conflicted misalignments in model predictions. Parallel pseudo-label ensembling reduces variability in pseudo-label selection by deriving consensus predictions from multiple OT-based models trained in parallel, thereby mitigating one-to-one misalignments. The ensembled pseudo-labeled alignments are then used to augment seed alignments to reinforce subsequent model training for alignment inference. To our best knowledge, we are the first to address the confirmation bias inherent in pseudo-labeling-based entity alignment. Comprehensive experiments and analyses validate the superior performance of UPL-EA over state-of-the-art supervised and semi-supervised baselines and its utility as a general pseudo-labeling framework to improve entity alignment performance.
@@ 2.1 Problem Statement
A knowledge graph (KG) can be represented as $\mathcal{G}=\{\mathcal{E},\mathcal{R},\mathcal{T}\}$ with the entity set $\mathcal{E}$, relation set $\mathcal{R}$, and relational triplet set $\mathcal{T}$. Each triplet is denoted as $(e_{i},r,e_{j})\in\mathcal{T}$, which represents that a head entity $e_{i}\in\mathcal{E}$ is connected to a tail entity $e_{j}\in\mathcal{E}$ via a relation $r\in\mathcal{R}$. Each entity $e_{i}$ is characterized by an entity feature vector $\bm{x}_{i}\in\mathbb{R}^{n}$, which can be obtained from entity names with semantic meanings.
Formally, given two KGs, $\mathcal{G}_{1}=\{\mathcal{E}_{1},\mathcal{R}_{1},\mathcal{T}_{1}\}$ and $\mathcal{G}_{2}=\{\mathcal{E}_{2},\mathcal{R}_{2},\mathcal{T}_{2}\}$, the task of entity alignment (EA) aims to discover a set of one-to-one equivalent entity pairs =$\mathcal{I}=\{(e_{i},e_{j})\in\mathcal{E}_{1}\times\mathcal{E}_{2}|\;e_{i}\equiv e_{j}\}$ between $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, where $e_{1}\in\mathcal{E}_{1}$, $e_{2}\in\mathcal{E}_{2}$, and $\equiv$ indicates an equivalence relationship between $e_{1}$ and $e_{2}$. In many cases, a small set of equivalent entity pairs $\mathcal{S}\subset\mathcal{I}$, known as prior seed alignments, is initially provided for training. Apart from the entities included in $\mathcal{S}$, there are two sets of unaligned entities $\mathcal{E}^{U}_{1}\subset\mathcal{E}_{1}$ and $\mathcal{E}^{U}_{2}\subset\mathcal{E}_{2}$ in $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$, respectively.
In this work, we address real-world scenarios where prior seed alignments $\mathcal{S}$ are often scarce due to high labeling costs. We focus on the task of pseudo-labeling-based entity alignment, which aims to leverage both prior seed alignments and unaligned entity pairs to more effectively train an EA model in a transductive semi-supervised setting. This is achieved by selecting a set of unaligned entity pairs as pseudo-labeled alignments $\mathcal{S}_{t}\subset\mathcal{E}^{U}_{1}\times\mathcal{E}^{U}_{2}$ in each iteration $t$, and using $\mathcal{S}_{t}$ to iteratively augment seed alignments $\mathcal{S}$, i.e., $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{t}$, for subsequent model training.
@@ 2.2 Analysis of Confirmation Bias
The key to pseudo-labeling-based entity alignment lies in selecting reliable pseudo-labeled alignments to effectively boost model performance; otherwise, pseudo-labeling errors could propagate into subsequent model training, leading to confirmation bias [1]. To investigate the impact of confirmation bias on pseudo-labeling-based entity alignment, we perform an error analysis of a naive pseudo-labeling strategy used in previous studies [41]. This strategy simply selects pairs of unaligned entities whose embedding distances (defined using Eq. (3)) are smaller than a pre-specified threshold as pseudo-labeled alignments. Our analysis is carried out on a widely used cross-lingual KG pair, DBP15KZH_EN, as detailed in Section 4.1. We follow the conventional 30%-70% split ratio to randomly partition 15,000 ground-truth alignments into training and test data. During model training, pseudo-labeled alignments are inferred from unaligned entity sets to augment seed alignments. To understand the underlying causes of confirmation bias, we explicitly calculate, in each pseudo-labeling iteration, the numbers of conflicted misalignments and one-to-one misalignments, as well as the number of correct one-to-one alignments, against ground-truth alignments on the test data.
As shown in Fig. 1(a), the naive strategy for selecting pseudo-labeled alignments introduces a substantial number of pseudo-labeling errors—including both conflicted misalignments and one-to-one misalignments—from the beginning. As training progresses, although the number of correct one-to-one alignments gradually increases owing to improved entity embeddings, both types of pseudo-labeling errors also accumulate and increase noticeably. This accumulation of errors gives rise to confirmation bias, severely hindering the capability of pseudo-labeling in improving entity alignment performance, as shown in Fig. 1(c).
(a) The naive pseudo-labeling strategy
(b) Our unified pseudo-labeling (UPL) strategy
(c) Entity alignment performance of two strategies
@@
Our analysis affirms that the confirmation bias essentially stems from pseudo-labeling errors. These errors, if not adequately addressed, would propagate through subsequent model training and jeopardize the effectiveness of pseudo-labeling for entity alignment. Motivated by this, our work focuses on explicitly identifying and eliminating the two types of pseudo-labeling errors: conflicted misalignments and one-to-one misalignments. As shown in Fig. 1(b), our proposed Unified Pseudo-Labeling (UPL) strategy can effectively eliminate all conflicted misalignments through OT-based pseudo-labeling and significantly reduce one-to-one misalignments via parallel pseudo-label ensembling across all iterations. Fig. 1(c) highlights the significant performance gains achieved by our proposed UPL strategy compared to the naive pseudo-labeling strategy.
@@ 3 The Proposed UPL-EA Framework
With insights from our analysis in Section 2.2, the proposed UPL-EA framework is designed to systematically address confirmation bias for pseudo-labeling-based entity alignment. The core of UPL-EA is a novel Unified Pseudo-Labeling (UPL) strategy that iteratively generates reliable pseudo-labeled alignments to enhance the training of entity alignment (EA) model. Essentially, UPL utilizes two key components: (1) OT-based pseudo-labeling, which generates pseudo-labeled alignments with one-to-one correspondences, effectively eliminating conflicted misalignments; and (2) parallel pseudo-label ensembling, which combines pseudo-labeled alignments generated from multiple OT-based models independently trained in parallel, reducing variability in pseudo-label selection and mitigating one-to-one misalignments.
Figure 2: An overview of the proposed UPL-EA framework. In pseudo-labeling iteration $t$, $M$ EA models are trained in parallel to generate a set of conflict-free pseudo-labeled alignments via OT-based pseudo-labeling. These alignments are further fed into parallel pseudo-label ensembling to generate ensembled pseudo-labeled alignments, which are then used to augment seed alignments for subsequent model training in the next iteration $t+1$.
Fig. 2 illustrates an overview of the proposed UPL-EA framework. In pseudo-labeling iteration $t$, an EA model learns entity embeddings based on a set of alignment seeds, $\mathcal{S}$, which are passed on to OT-based pseudo-labeling to generate conflict-free pseudo-labeled alignments. Rather than relying on predictions from a single model that are potentially unreliable, $M$ EA models are independently trained in parallel to generate a set of conflict-free pseudo-labeled alignments $\{\widehat{\mathcal{S}}_{t}^{(1)},\widehat{\mathcal{S}}_{t}^{(2)},...,\widehat{\mathcal{S}}_{t}^{(M)}\}$ through OT-based pseudo-labeling. These alignments are then combined through parallel pseudo-label ensembling, which retains only the consensus pseudo-labeled alignments generated from $M$ models, i.e., $\mathcal{S}_{t}=\cap_{m=1}^{M}\widehat{\mathcal{S}}_{t}^{(m)}$. The ensembled pseudo-labeled alignments $\mathcal{S}_{t}$ are then used to augment seed alignments, i.e., $\mathcal{S}\leftarrow\mathcal{S}\cup\mathcal{S}_{t}$, for subsequent model training in the next pseudo-labeling iteration $t+1$. Through this iterative process, the EA models and the UPL strategy mutually reinforce each other, progressively leading to more informative entity embeddings. Finally, the learned entity embeddings are used for entity alignment inference.